OpenStudy (anonymous):

Differentiate (2x)/(1-3x^2)

OpenStudy (anonymous):

quotient rule?

OpenStudy (anonymous):

i guess, thats all the problem says

OpenStudy (anonymous):

denominator will be \[(1-3x^2)^2\] that is the easy part

OpenStudy (anonymous):

numerator will be \[ (1-3x^2)\times 2-2x\times (-6x)\] then some algebra

OpenStudy (anonymous):

no that is wrong, i get \[\frac{6x^2-2}{(1-3x^2)^2}\]

OpenStudy (anonymous):

thanks, that is what i got too...just wanted to be sure, he never worked one like that

OpenStudy (anonymous):

yw